8) (5 pts) Light with wavelength of 650 nm shines down on a curved piece of glass that rests on a flat glass as shown. The boPlease explain how to solve this. The answer should be B
Answer #1

The effective path difference between two reflected Jays for a wedge shaped film 4 = 2 Mte cos (3+6) + Refractive index ty Th.. . 2 + + = А 2 25 = (2n-1) 2. 4= t2-1) А м= 1, 2-,.. 41 - О not allowed :. А= (2xi - 1) А 4 Бол ? 1) н 650 mm Ах1.3 125 mm